OpenStudy (anonymous):

Can someone give me the value of x? 5.6/8=X/24

OpenStudy (shamim):

8x=24*5.6 x=?

OpenStudy (anonymous):

24*5.6= 134.4

OpenStudy (♪chibiterasu):

Now divide 8 on both sides.

OpenStudy (anonymous):

16.8 is the answer right?

OpenStudy (shamim):

Ya ur answer is correct!!!